Solve by completing the square
x^2+9x+20=0 and can you show me how you got that

Answer by scott8148(6628)

subtracting 20 ___ x^2 + 9x = -20

when the coefficient of the squared term is one (as in this case)
___ take half of the 1st order coefficient , square the quantity , add to both sides of equation

x^2 + 9x + (9/2)^2 = -20 + (9/2)^2

[x + (9/2)]^2 = -(80/4) + (81/4)

x + (9/2) = sqrt(1/4)

x = -(9/2) ± (1/2)

x = -5 ___ and ___ x = -4